OpenStudy (anonymous):

−2(x + 3) = −2(x + 1) − 4

OpenStudy (bahrom7893):

-2x-6=-2x-2-4

OpenStudy (bahrom7893):

True for all x

OpenStudy (anonymous):

Distribute first. -2(x + 3) = -2(x + 1) - 4 -2x -6 = -2x - 1 - 4 Combine like terms. Then solve for x. Does that help?

OpenStudy (anonymous):

yes

OpenStudy (bahrom7893):

@jazy, wrong -2(x + 1) - 4 = -2x-2-4

OpenStudy (anonymous):

this a two step

OpenStudy (anonymous):

@bahrom7893 that's what I meant to type in. So: -2x -6 = -2x - 2 - 4